In a study of 784 randomly selected medical malpractice? lawsuits, it was found that 492 of them were dropped or dismissed. Use a 0.05 significance level to test the claim that most medical malpractice lawsuits are dropped or dismissed.
Which of the following is the hypothesis test to be? conducted?
A.
Upper H 0 : p less than 0.5
Upper H 1 : p equals 0.5
B.
Upper H 0 : p greater than 0.5
Upper H 1 : p equals 0.5
C.
Upper H 0 : p equals 0.5
Upper H 1 : p not equals 0.5
D.
Upper H 0 : p equals 0.5
Upper H 1 : p less than 0.5
E.
Upper H 0 : p not equals 0.5
Upper H 1 : p equals 0.5
F.
Upper H 0 : p equals 0.5
Upper H 1 : p greater than 0.5


What is the test? statistic?
Z =
?(Round to two decimal places as? needed.)


What is the conclusion about the null? hypothesis?
A. Reject the null hypothesis because the? P-value is greater than the significance? level, alpha.
B. Fail to reject the null hypothesis because the? P-value is less than or equal to the significance? level, alpha.
C.Fail to reject the null hypothesis because the? P-value is greater than the significance? level, alpha.
D.Reject the null hypothesis because the? P-value is less than or equal to the significance? level, alpha.
What is the final? conclusion?
A.There is sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.
B.There is not sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.
C.There is sufficient evidence to warrant rejection of the claim that most medical malpractice lawsuits are dropped or dismissed.
D.There is not sufficient evidence to warrant rejection of the claim that most medical malpractice lawsuits are dropped or dismissed.
Answer #1

Ans:

Option F is correct.

H0:p=0.5

H1:p>0.5

Sample proportion=492/784=0.6276

Test statistic:

z=(0.6276-0.5)/sqrt(0.5*(1-0.5)/784)

z=7.14

p-value=P(z>7.14)=0

As,p-value<0.05,we reject null hypothesis.

Option D is correct.(.Reject the null hypothesis because the? P-value is less than or equal to the significance? level, alpha.)

Option A is correct.(There is sufficient evidence to support the claim that most medical malpractice lawsuits are dropped or dismissed.)